# PixelCAM: Pixel Class Activation Mapping for Histology Image Classification and ROI Localization

### v1

@@ Abstract
Weakly supervised object localization (WSOL) methods allow training models to classify images and localize ROIs. WSOL only requires low-cost image-class annotations yet provides a visually interpretable classifier, which is important in histology image analysis. Standard WSOL methods rely on class activation mapping (CAM) methods to produce spatial localization maps according to a single- or two-step strategy. While both strategies have made significant progress, they still face several limitations with histology images. Single-step methods can easily res4ult in under- or over-activation due to the limited visual ROI saliency in histology images and the limited localization cues. They also face the well-known issue of asynchronous convergence between classification and localization tasks. The two-step approach is sub-optimal because it is tied to a frozen classifier, limiting the capacity for localization. Moreover, these methods also struggle when applied to out-of-distribution (OOD) datasets. In this paper, a multi-task approach for WSOL is introduced for simultaneous training of both tasks to address the asynchronous convergence problem. In particular, localization is performed in the pixel-feature space of an image encoder that is shared with classification. This allows learning discriminant features and accurate delineation of foreground/background regions to support ROI localization and image classification. We propose PixelCAM, a cost-effective foreground/background pixel-wise classifier in the pixel-feature space that allows for spatial object localization. Using partial-cross entropy, PixelCAM is trained using pixel pseudo-labels collected from a pretrained WSOL model. Both image and pixel-wise classifiers are trained simultaneously using standard gradient descent. In addition, our pixel classifier can easily be integrated into CNN- and transformer-based architectures without any modifications. Our extensive experiments [^1] Our public code: github.com/AlexisGuichemerreCode/PixelCAM on GlaS and CAMELYON16 cancer datasets show that PixelCAM can improve classification and localization performance when integrated with different WSOL methods. Most importantly, it provides robustness on both tasks for OOD data linked to different cancer types, with large domain shifts between training and testing image data.
Keywords: Deep Learning, Image Classification, Visual Interpretability, Weakly Supervised Object Localization, Histology Images, Out-Of-Distribution Data
@@ 1 Introduction
Histology image analysis remains the gold standard for diagnosing cancers of the brain [15], breast [25], and colon [28]. However, training deep models for accurate localization of cancerous regions of interest (ROIs) requires pixel-wise annotation by pathologists which is a complex and time-consuming task, especially over Whole Slide Images (WSI). Weakly supervised object localization (WSOL) has emerged as a low-cost training approach [32]. Using only image class supervision, a deep model can be trained to classify an image, but also to provide spatial localization of ROIs associated with that class [3, 7, 17, 22]. This significantly reduces the large cost of annotation and the need for dense pixel supervision. In addition, it builds interpretable classifiers [4, 19] which is critical in medical domains such as in histology image analysis.
@@
Recent progress in WSOL is dominated by class activation mapping (CAM) methods [22]. Several methods extract spatial activation maps in a single step [13, 20, 27], and require a module on top of a feature extractor to construct these maps. Then, a spatial pooling module extracts per-class scores used for supervised training using image-class labels. However, without any explicit guidance at the pixel level, these CAM methods can lead to poor maps due to the challenging nature of histology images where objects are often not salient [22]. This can lead to under- or over-activation which creates a high false negative/positive rate at the pixel level. Moreover, training a single model may face the issue of asynchronous convergence of classification and localization tasks [7, 17, 22].
@@
To bypass this convergence issue and the lack of pixel guidance, a recent direction has emerged in WSOL where it aims at providing explicit localization cues as pseudo-labels with dual models [3, 5, 18, 26, 29, 30]. In particular, a per-task model is considered where the localizer is trained using pseudo-labels [30]. This leads to more parameters and training cycles. In addition, both tasks are disconnected leading to unrelated decisions in both models. A different approach [3] uses a single model where a decoder is combined with a frozen classifier. While this yields good results, this architecture is limited as it is tied to frozen features at many layers using skip connections. This prevents the localizer from a better adaptation and hinders its performance.
@@
In addition, recent work has revealed a major limitation to WSOL methods when dealing with domain shift in histology analysis [12]. The performance of WSOL methods is shown to decline on both tasks with out-of-distribution (OOD), limiting their real-world application. Further inspection suggests that features at the pixel level can be the root cause. Since they lack direct localization supervision, a feature encoder may provide poorly separated features concerning classes, making them vulnerable to domain shift, and class confusion.
@@
To alleviate the aforementioned issues, we propose a novel multi-task WSOL method called PixelCAM for histology image analysis. It is based on a cost-effective foreground (FG)/background (BG) pixel-wise classifier working in the pixel-feature space, allowing for spatial object localization. It aims to explicitly learn discriminant pixel-feature representations and accurate delineation of FG and BG regions, to provide better ROI localization and image classification. This training is achieved through localization pseudo-labels extracted from a pretrained WSOL model. In addition, such separation provides the model with robustness to OOD data. This creates a single-step multi-task framework that performs simultaneously classification and localization tasks. Multi-task training is therefore used for a better adaptation by learning rich features for both tasks, compared to tied features in a two-step approach which limits learning. This also makes both tasks cooperate to converge to a satisfying solution for the two tasks. The multi-task optimization is done using standard gradient descent by using provided image-class supervision and pixel-wise pseudo-labels extracted from a pretrained WSOL CAM model.
Our main contributions are summarized as follows.
@@
(1) A novel multi-task WSOL method called PixelCAM is proposed for histology image analysis. Multi-task optimization of both classification and localization tasks is achieved in a single step by integrating a pixel classifier at the pixel-feature level while leveraging localization pseudo-labels. PixelCAM alleviates the asynchronous convergence issue in WSOL, improves the performance of both tasks, and importantly, provides robustness to OOD data. Our pixel classifier is versatile in that it can easily be integrated into any CNN- or transform-based classifier architecture without modification.
(2) We conduct extensive experiments on two public datasets for colon (GlaS), and breast cancer (CAMELYON16). Our method outperforms WSOL baseline methods on both tasks. Additionally, when dealing with large domain shifts across cancer types, our PixelCAM method can maintain a high level of accuracy, making it an ideal choice for OOD scenarios. We provide several ablations to further analyze our method.
Figure 1: Top row: WSOL learning strategies. (a) Single-step WSOL methods perform classification and localization in one step. While standard WSOL uses a sequential approach using only image-class and lacks leveraging localization cues, PixelCAM relies on a multi-task training using both image-class and localization cues. (b) Two-step WSOL methods use a model per task or an encoder-decoder which are trained sequentially.
Bottom row: PixelCAM training (c-1) and inference (c-2). The training relies on pseudo-labels collected from a WSOL CAM classifier pretrained on the same dataset ${\mathbb{D}}$. The classification head aims to classify the extracted features ${\bm{F}}$. The pixel-wise classifier predicts the randomly selected locations as foreground or background. At test time, all locations are classified using the localization head ${f}$ producing a localization map, while classifying the whole image using the classification head ${g}$.
2 Related Work
WSOL has emerged as a low-cost training setup to localize objects while classifying an image content [7, 17, 32]. WSOL has also been extended to videos [1, 2]. Moreover, there has been a recent focus on designing WSOL methods for histology image analysis [22]. This allows us to reduce the large annotation cost, in addition to building visually interpretable classifiers.
Single-step WSOL. Several approaches address the WSOL problem in a single step [7, 17, 22] where a single model is trained to do both tasks, classification and localization. Usually, this is achieved by using a localization head followed by a spatial pooling layer to extract per-class probability. Early [13, 23, 31] but also recent WSOL works [27, 33, 11] follow this strategy. Although this has brought significant improvements into the field, these models still face several limitations. Since only image-class labels are used as supervision without any localization cues, CAM localization can have poor performance, more so when dealing with less salient objects such as in histology images. Recent work [22] showed that without localization cues over histology images, these models often lead to highly unbalanced localization. This manifests either by under- or over-activations leading to high false negative/positive rates. In addition, this approach faces a great challenge in model selection as both tasks proceed in an asynchronous convergence [7, 17, 22]. For instance, the model selected for best localization often yields poor classification.
Two-step WSOL.This strategy has emerged as a new line of research. In particular, a dual-model approach is used in combination with the usage of localization cues under the form of pseudo-labels [17, 22]. This provides direct guidance for the localization task, bypassing the issue of task convergence. Several works simply create a dedicated model per task: one for classification and another for localization [26, 29, 30]. This leads to better performance since tasks are divided between two large models. However, this is a cumbersome approach as it drastically increases the number of parameters and training cycles. Most importantly, the localization CAMs are completely disconnected from the classification decision, making this strategy unreliable and less interpretable. A parallel direction overcomes this issue by using a decoder to act as a localizer on top of the classifier [3, 5, 18], in a similar way to a U-Net architecture [21]. This creates a direct relation between both tasks. However, the decoder has a limited learning capacity since it is tied to a frozen encoder at several layers via skip connections. This prevents the localizer from a better adaptation, and it limits its performance leading to a sub-optimal solution. In addition, it adds a significant number of parameters to the classifier.
WSOL vs OOD. In addition to these limitations, WSOL methods face challenges when dealing with domain shift which degrades the performance of both tasks [12]. Such shift is common in histology due to variations in stains, objects’ structure, microscope type, and imaging centers. Further analysis shows that this issue could be rooted in the pixel features of the image encoder. Both classification and localization tasks depend heavily on these spatial features. Less flexible and poorly discriminant pixel features can lead to poor performance in both tasks since they both rely on these embeddings.
@@ 3 Proposed Method
Let us denote by ${\mathbb{D}=\{(\bm{X},y)_{i}\}_{i=1}^{N}}$ a training set with $N$ samples, where ${\bm{X}:\Omega\subset\mathbb{R}^{2}}$ is an input image and ${y\in\{1,\cdots,K\}}$ is its global label, with $K$ being the number of classes. Our model is composed of three parts (Fig.1): (a) feature extractor, (b) image classifier, and (c) pixel-wise classifier. (a) The feature extractor backbone $h$ with parameter ${\bm{\theta}_{1}}$ produces a tensor spatial features ${h(\bm{X})=\bm{F}\in\mathbb{R}^{\left|\Omega\right|\times d}}$, with depth $d$. For simplicity, we consider that $\bm{F}$ has the same dimensions as the input image ${\bm{X}}$ through interpolation. (b) The global image classifier head $g$ classifies the image content, with parameters ${\bm{\theta}_{2}}$, where ${g(\bm{F})}$ is the per-class probability. (c) The pixel-wise [^2] For simplicity, we refer to a location in a spatial tensor as a pixel. However, in DL models, such location typically covers more than one pixel in the image space due to the receptive field. Interpolation can be used to upscale the features to the full image size. classifier $f$ which classifies the embedding ${\bm{F}_{p}\in\mathbb{R}^{d}}$ of a pixel at location $p$ in ${\bm{F}}$ with parameters ${\bm{\theta}_{3}}$ into either foreground (1) or background (0) classes. This is typically a linear classifier. We refer by ${f(\bm{F}_{p})_{0}}$, ${f(\bm{F}_{p})_{1}}$ as the probability for the pixel at location $p$ to be background and foreground, respectively. For simplicity, we use ${f(\bm{F}_{p})}$ for pixel-class probability. Classifying all locations creates the two localization maps ${\bm{S}=f(\bm{F})\in[0,1]^{\left|\Omega\right|\times 2}}$, where ${\bm{S}^{0},\bm{S}^{1}}$ refer to the background and foreground maps, respectively. For simplicity, we refer by ${y^{\prime}\in\{0,1\}}$ as the pseudo-label of a pixel location ${p}$. Let us denote the standard cross-entropy by $\bm{H}(\cdot,\cdot)$. The total set of parameters of our model is referred to as ${\bm{\theta}=\{\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{3}\}}$. When ${y,y^{\prime}}$ are used with ${\bm{H}}$, it is the one-hot encoding version that is being considered.
@@
$\min_{\bm{\theta}_{1},\bm{\theta}_{2}}$ $\bm{H}(y,g(\bm{F}))\;.$ (1)
@@
Pixel-wise classifier. Training the model using only Eq.1 builds a spatial feature ${\bm{F}}$ that is guided only by the image classification loss. It therefore lacks awareness of localization as it does not have access to direct localization supervision cues. This can easily lead to poor pixel-wise features and confuse classes at the feature level. This is particularly true when dealing with less salient and strongly similar objects, as found in histology images. Subsequently, localization on top of these features is less reliable. Most importantly, introducing domain shift over the input image will lead to further feature confusion and failure of subsequent modules to localize and classify. To mitigate this issue, our pixel-classifier $f$ is used to create well-separated features between foreground/background in the pixel-feature space.
@@
Training $f$ requires localization cues. Since the WSOL setup does not provide them, we resort to using a pretrained WSOL CAM classifier trained on the same train dataset ${\mathbb{D}}$, based on CNN or transformer, to acquire pseudo-labels. This approach is typically employed in WSOL to train a new DL model or a decoder for localization [22]. However, this is cumbersome and requires another training phase with many additional parameters. To prevent this issue, our method trains a linear model to classify pixel embeddings while training the image classifier and sharing a full backbone. The number of parameters and training cycles is thereby reduced considerably.
@@
To produce pixel-wise pseudo labels, standard WSOL methods [22] are adopted. We use the CAM ${\bm{C}}$ of the true image class $y$ from a WSOL CAM classifier pretrained on the same train dataset ${\mathbb{D}}$. Strong activations typically indicate foreground regions, while low activates point to background regions [9]. Instead of directly fitting these regions, recent works showed that it is better to stochastically sample regions to avoid overfitting [18, 5]. To sample $n$ random foreground pixel locations, a multinomial distribution is used with ${\bm{C}}$ as its pixel-sampling probability. Typically, this allows for sampling more frequently from high-activation locations. For background sampling, ${1-\bm{C}}$ is instead used as a sampling probability. Sampled pixels are collected in the sets ${\mathbb{C}^{+},\mathbb{C}^{-}}$ for foreground and background pixels, respectively. Note that sampled locations change for each image at every training step. This allows to explore different regions, and it reduces overfitting to specific regions. The sampled pseudo-labels can be directly used to train our pixel classifier, using partial cross-entropy since only part of the image space ${\Omega}$ is considered at once,
$\min_{\bm{\theta}_{1},\bm{\theta}_{3}}$ $\sum_{p\in\{\mathbb{C}^{+}\;\cup\;\mathbb{C}^{-}\}}\bm{H}(y^{\prime},f(\bm{F}_{p}))\;.$ (2)
@@
$\min_{\bm{\theta}}$ $\bm{H}(y,g(\bm{F}))+\lambda\sum_{p\in\{\mathbb{C}^{+}\;\cup\;\mathbb{C}^{-}\}}\bm{H}(y^{\prime},f(\bm{F}_{p}))\;,$ (3)
where ${\lambda}$ is a balancing coefficient. Eq.3 is minimized using standard gradient descent to perform both tasks. This mitigates the convergence issue, and allows a single training cycle. It also ensures a cooperation between localization and classification to reach a better solution for both tasks at the same time. PixelCAM does not involve a considerable computation overhead. Given our lightweight pixel classifier, the number of additional parameters is negligible compared to standard DL models. Moreover, this pixel classifier is generic and can be integrated into CNN- or transformer-based models without architectural changes. Our pixel-classifier creates a boundary in the pixel feature space to well separate ROIs from noise and increase the feature’s discriminant power. This helps localization, but also the subsequent classification module as they use the same spatial features for decision. Consequently, it makes features robust to OOD data, allowing better performance of both tasks in such scenario. In Section 4, we show that including this linear classifier into a standard deep WSOL model, allows us to improve its localization and classification accuracy. Moreover, it provides robustness to domain shift, a common issue in WSOL models for histology image analysis [12].

### v2

@@ Abstract
[…]
keywords:
Deep Learning, Image Classification, Visual Interpretability, Weakly Supervised Object Localization, Histology Images, Out-Of-Distribution Data.
editors: Accepted for publication at MIDL
[…]
@@ 1 Introduction
Histology image analysis remains the gold standard for diagnosing cancers of the brain [Khalsa et al. (2020)], breast [Veta et al. (2014)], and colon [Xu et al. (2020)]. However, training deep learning (DL) models for accurate localization of cancerous regions of interest (ROIs) requires pixel-wise annotation by pathologists which is a complex and time-consuming task, especially over Whole Slide Images (WSI). WSOL has emerged as a low-cost training approach [Zhou (2017)]. Using only image class supervision, a DL model can be trained to classify an image, but also to provide spatial localization of ROIs associated with that class [Belharbi et al. (2022a), Choe et al. (2020), Murtaza et al. (2024), Rony et al. (2023)]. This significantly reduces the large cost of annotation and the need for dense pixel supervision. In addition, it builds interpretable classifiers [Belharbi et al. (2022b), Neto et al. (2024)] which is critical in medical domains such as in histology image analysis.
@@
Recent progress in WSOL is dominated by CAM methods [Rony et al. (2023)]. Several methods extract spatial activation maps in a single step [Ilse et al. (2018), Oquab et al. (2015), Wu et al. (2023)], and require a module on top of a feature extractor to construct these maps. Then, a spatial pooling module extracts per-class scores used for supervised training using image-class labels. However, without any explicit guidance at the pixel level, these CAM methods can lead to poor maps due to the challenging nature of histology images where objects are often not salient [Rony et al. (2023)]. This can lead to under- or over-activation which creates a high false negative/positive rate at the pixel level. Moreover, training a single model may face the issue of asynchronous convergence of classification and localization tasks [Choe et al. (2020), Murtaza et al. (2024), Rony et al. (2023)].
@@
To bypass this convergence issue and the lack of pixel guidance, a recent direction has emerged in WSOL where it aims at providing explicit localization cues as pseudo-labels with dual models [Belharbi et al. (2022a), Belharbi et al. (2022c), Murtaza et al. (2023), Wei et al. (2021), Zhang et al. (2020), Murtaza et al. (2025), Murtaza et al. (2022b), Zhao et al. (2023), Murtaza et al. (2022a)]. In particular, a per-task model is considered where the localizer is trained using pseudo-labels [Zhao et al. (2023)]. This leads to more parameters and training cycles. In addition, both tasks are disconnected leading to unrelated decisions in both models. A different approach [Belharbi et al. (2022a)] uses a single model where a decoder is combined with a frozen classifier. While this yields good results, this architecture is limited as it is tied to frozen features at many layers using skip connections. This prevents the localizer from a better adaptation and hinders its performance.
@@
In addition, recent work has revealed a major limitation to WSOL methods when dealing with domain shift in histology analysis [Guichemerre et al. (2024)]. The performance of WSOL methods is shown to decline on both tasks with out-of-distribution (OOD), limiting their real-world application. Further inspection suggests that features at the pixel level can be the root cause. Since they lack direct localization supervision, a feature encoder may provide poorly separated features concerning classes, making them vulnerable to domain shift, and class confusion.
@@
To alleviate the aforementioned issues, we propose a novel multi-task WSOL method for histology image analysis. It is based on called PixelCAM– a cost-effective foreground / background (FG/BG) pixel-wise classifier working in the pixel-feature space, allowing for spatial object localization. It aims to explicitly learn discriminant pixel-feature representations and accurate delineation of FG and BG regions. This improves the ROI localization and image classification accuracy. Training is achieved through localization pseudo-labels extracted from a pretrained WSOL model. In addition, such pixel-wise classification provides the model with robustness to OOD data. Our single-step multi-task framework allows to simultaneously perform classification and localization tasks. Multi-task training is therefore leveraged to learn rich features for both tasks, compared to the constrained learning of features in a two-step approach. Our approach cooperates to converge to a satisfying solution for both tasks. The multi-task optimization is performed using standard gradient descent by using image-class labels and pixel-wise pseudo-labels extracted from a pretrained WSOL model.
@@
Our main contributions are summarized as follows. (1) A novel multi-task WSOL method called PixelCAM is proposed for histology image analysis. Multi-task optimization of both classification and localization tasks is achieved in a single step by integrating a pixel classifier at the pixel-feature level while leveraging localization pseudo-labels. PixelCAM alleviates the asynchronous convergence issue in WSOL, improves the performance of both tasks, and importantly, provides robustness to OOD data. Our pixel classifier is versatile in that it can easily be integrated into any CNN- or transform-based classifier architecture without modification. (2) We conduct extensive experiments on two public datasets for colon (GlaS), and breast cancer (CAMELYON16). Our method outperforms WSOL baseline methods on both tasks. Additionally, when dealing with large domain shifts across cancer types, our PixelCAM method can maintain a high level of accuracy, making it an ideal choice for OOD scenarios. We provide several ablations to further analyze our method.
@@ 2 Proposed Method
Figure 1: Top row: WSOL learning strategies. (a) Single-step WSOL methods perform classification and localization in one step. While standard WSOL uses a sequential approach using only image-class and lacks leveraging localization cues, PixelCAM relies on a multi-task training using both image-class and localization cues. (b) Two-step WSOL methods use a model per task or an encoder-decoder which are trained sequentially. Bottom row: PixelCAM training (c-1) and inference (c-2). The training relies on pseudo-labels collected from a WSOL CAM classifier pretrained on the same dataset ${\mathbb{D}}$. The classification head aims to classify the extracted features ${\mathsf{F}}$. The pixel-wise classifier predicts the randomly selected locations as foreground or background. At test time, all locations are classified using the localization head ${f}$ producing a localization map, while classifying the whole image using the classification head ${g}$.
Let us denote by ${\mathbb{D}=\{(\bm{X},y)_{i}\}_{i=1}^{N}}$ a training set with $N$ samples, where ${\bm{X}:\Omega\subset\mathbb{R}^{2}}$ is a 2d input image of dimension ${h^{\prime}\times w^{\prime}}$ and ${y\in\{1,\cdots,K\}}$ is its global label, with $K$ being the number of classes. Our model is composed of three parts (Fig.1): (a) feature extractor, (b) image classifier, and (c) pixel-wise classifier. (a) The feature extractor backbone $h$ with parameter ${\bm{\theta}_{1}}$ produces a tensor spatial features ${h(\bm{X})=\mathsf{F}\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times d}}$, with depth $d$. For simplicity, we consider that $\mathsf{F}$ has the same dimensions as the input image ${\bm{X}}$ through interpolation. (b) The global image classifier head $g$ classifies the image content, with parameters ${\bm{\theta}_{2}}$, where ${g(\mathsf{F})}$ is the per-class probability. (c) The pixel-wise [^2] For simplicity, we refer to a location in a spatial tensor as a pixel. However, in DL models, such location typically covers more than one pixel in the image space due to the receptive field. Interpolation can be used to upscale the features to the full image size. classifier $f$ which classifies the embedding ${\mathsf{F}_{i,j,:}=\mathsf{F}_{p}\in\mathbb{R}^{d}}$ of a pixel at location ${(i,j,:)}$ or simply $p$ in ${\mathsf{F}}$ with parameters ${\bm{\theta}_{3}}$ into either foreground (1) or background (0) classes. This is typically a linear classifier. We refer by ${f(\mathsf{F}_{p})_{0}}$, ${f(\mathsf{F}_{p})_{1}}$ as the probability for the pixel at location $p$ to be background and foreground, respectively. For simplicity, we use ${f(\mathsf{F}_{p})}$ for pixel-class probability. Classifying all locations creates the two localization maps ${\mathsf{S}=f(\mathsf{F})\in[0,1]^{h^{\prime}\times w^{\prime}\times 2}}$, where ${\mathsf{S}_{:,:,0},\mathsf{S}_{:,:,1}}$ or simply ${\mathsf{S}^{0},\mathsf{S}^{1}}$ refer to the background and foreground maps, respectively. For simplicity, we refer by ${y^{\prime}\in\{0,1\}}$ as the pseudo-label of a pixel location ${p}$. Let us denote the standard cross-entropy by $\mbox{H}(\cdot,\cdot)$. The total parameters of our model is referred to as ${\bm{\theta}}$ which is composed of ${\{\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{3}\}}$. When ${y,y^{\prime}}$ are used with H, it is the one-hot encoding version that is being considered.
@@
$\min_{\bm{\theta}_{1},\bm{\theta}_{2}}$ $\mbox{H}(y,g(\mathsf{F}))\;.$ (1)
@@
Pixel-wise classifier. Training the model using only Eq.1 builds a spatial feature ${\mathsf{F}}$ that is guided only by the image classification loss. It therefore lacks awareness of localization as it cannot directly access localization supervision cues. This may lead to poor pixel-wise features and confuse classes at the feature level. This is particularly true when dealing with less salient and strongly similar objects, as found in histology images. Subsequently, localization on top of these features is less reliable. Most importantly, introducing domain shift over the input image will lead to further feature confusion and failure by subsequent modules to localize and classify. To mitigate this issue, our pixel-classifier $f$ is used to create well-separated features between foreground/background in the pixel-feature space.
@@
Training $f$ requires localization cues. Since the WSOL setup does not provide them, we resort to using a pretrained WSOL CAM classifier trained on the same train dataset ${\mathbb{D}}$, based on CNN or transformer, to acquire pseudo-labels. This approach is typically employed in WSOL to train a new DL model or a decoder for localization [Rony et al. (2023)]. However, this is cumbersome and requires another training phase with many additional parameters. To prevent this issue, our method trains a linear model to classify pixel embeddings while training the image classifier and sharing a full backbone. The number of parameters and training cycles is thereby reduced considerably.
@@
To produce pixel-wise pseudo labels, standard WSOL methods (Rony et al., 2023) are adopted. We use the CAM ${\bm{C}}$ of the true image class $y$ from a WSOL CAM classifier pretrained on the same train dataset ${\mathbb{D}}$. Strong activations typically indicate foreground regions, while low activates point to background regions Durand et al. (2017). Instead of directly fitting these regions, recent works showed that it is better to stochastically sample regions to avoid overfitting Murtaza et al. (2023); Belharbi et al. (2022c); Murtaza et al. (2022a). To sample $n$ random foreground pixel locations, a multinomial distribution is used with ${\bm{C}}$ as its pixel-sampling probability. Typically, this allows for sampling more frequently from high-activation locations. For background sampling, ${1-\bm{C}}$ is instead used as a sampling probability. Sampled pixels are collected in the sets ${\mathbb{C}^{+},\mathbb{C}^{-}}$ for foreground and background pixels, respectively. Note that sampled locations change for each image at every training step. This allows to explore different regions, and it reduces overfitting to specific regions. The sampled pseudo-labels can be directly used to train our pixel classifier, using partial cross-entropy since only part of the image space ${\Omega}$ is considered at once,
$\min_{\bm{\theta}_{1},\bm{\theta}_{3}}$ $\sum_{p\in\{\mathbb{C}^{+}\;\cup\;\mathbb{C}^{-}\}}\mbox{H}(y^{\prime},f(\mathsf{F}_{p}))\;.$ (2)
@@
$\min_{\bm{\theta}}$ $\mbox{H}(y,g(\mathsf{F}))+\lambda\sum_{p\in\{\mathbb{C}^{+}\;\cup\;\mathbb{C}^{-}\}}\mbox{H}(y^{\prime},f(\mathsf{F}_{p}))\;,$ (3)
where ${\lambda}$ is a balancing coefficient. Eq.3 is minimized using standard gradient descent to perform both tasks. This mitigates the convergence issue, and allows a single training cycle. It also ensures a cooperation between localization and classification to reach a better solution for both tasks at the same time. PixelCAM does not involve a considerable computation overhead. Given our lightweight pixel classifier, the number of additional parameters is negligible compared to standard DL models. Moreover, this pixel classifier is generic and can be integrated into CNN- or transformer-based models without architectural changes. Our pixel-classifier creates a boundary in the pixel feature space to well separate ROIs from noise and increase the feature’s discriminant power. This helps localization, but also the subsequent classification module as they use the same spatial features for decision. Consequently, it makes features robust to OOD data, allowing better performance of both tasks in such scenario. In Section 3, we show that including this linear classifier into a standard deep WSOL model, allows us to improve its localization and classification accuracy. Moreover, it provides robustness to domain shift, a common issue in WSOL models for histology image analysis Guichemerre et al. (2024).